OpenStudy (anonymous):

the solution set sqrt3x+16=x+2 is (only 3x+16 is squared)

OpenStudy (anonymous):

\[\sqrt{3x+16}=x+2\]??

OpenStudy (anonymous):

square both sides get \[3x+16=x^2+4x+4\] then solve \[x^2+x-12=0\] \[(x-3)(x+4)=0\] \[x=3\] or \[x=-4\] then check the answers because only one will work